The Garraty Company has two bond issues outstanding. Both bonds pay $100 annual interest plus $1,000 at maturity. Bond L has a maturity of 15 years, and Bond S has a maturity of 1 year. a. 1. what will be the value of each of these bonds when the going rate of interest is 4%? Assume that there is only one more interest payment to be made on Bond S. Round your answers to the nearest cent. Bond L:$ Bond S: 2, what will be the value of each of these bonds when the going rate of interest is 7%? Assume that there is only one more interest payment to be made on Bond S. Round your answers to the nearest cent. Bond L: Bond S: $ 3, what will be the value of each o these bonds when the going rate of interest is 12%? Assume that there is on / one more interest paymentto e made on Bonds. Round you answers the nearest cent. Bond L: 863.78 Bond S: | 982.14 b. Why does the longer-term (15-year) bond fluctuate more when interest rates change than does the shorter-term bond (1 year)? I. Shorter-term bonds have more interest rate risk than longer-term bonds. II. Longer-term bonds have more interest rate risk than shorter-term bonds. III. Longer-term bonds have more reinvestment rate risk than shorter-tm bonds -Select- ▼
Answer #1

a.

1. Value of Bond L is $100 * 11.11839 + $1,000 * 0.55527 = $1,667.11

Value of Bond S is $100 * 0.96154 + $1,000 * 0.96154 = $1,057.69

2. Value of Bond L is $100 * 9.10791 + $1,000 * 0.36245 = $1,273.24

Value of Bond S is $100 * 0.93458 + $1,000 * 0.93458 = $1,028.04

b. Longer term bond fluctuate more when interest rate changes than does the shorter term bond because Longer-term bonds have more reinvestment rate risk than short-term bonds i.e. option III.